40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4523     Entry time: Thu Apr 14 01:03:43 2011
Author: Koji 
Type: Update 
Category: LSC 
Subject: LSC Campaign ~ Status 

Target: To lock the Michelson with the new RF/LSC


Status

RF generation box: READY - already ready to go to the IOO rack. (Suresh)

RF distribution box: In Progress - the internal components are to be connected. (13th evening - Suresh)

Placing PD and CCD: Done - PD and CCD on the AP table (13th Afternoon - Aidan, Larisa with supervision of Kiwamu)

Cabling1: Done - PD signal AP table to the demodulator (13th Afternoon - Jamie with supervision of Suresh)

Cabling2: Done - RF generation box (IOO Rack) to the demodulator

Demodulator: In Progress - Test and install (13th night - Kiwamu with supervision of Suresh)

LSC model: Done - Run the new LSC model. (It is named as "C1LST" so far) (13th evening - Jamie)

LSC medm: Done on 14th - Modify the current LSC medm screens Update the EPICS database Adjust the matrices (- Jenne with supervision of Koji)

ELOG V3.1.3-